OpenStudy (goformit100):

If 1+2=5 3+4 =19 5+6=41 then 1+3=?

OpenStudy (hexagon001):

10

OpenStudy (goformit100):

@uri Help me Madam.

OpenStudy (hexagon001):

the second term in each of the expressions is squared then added to the first. for example 1+2=4 the second term is 2... square this... this gives 4... now add it to the first term, which is 1. so 4+1=5

OpenStudy (perl):

1 + 2 is not equal to 5

OpenStudy (hexagon001):

@perl this is a sequence of numbers and you must figure how it is so.. see my post above for explanation

OpenStudy (hexagon001):

@goformit100 do you get it?

OpenStudy (goformit100):

Please give the detailed explanation

OpenStudy (hexagon001):

i have already done so ...five posts up

OpenStudy (hexagon001):

@goformit100 if you still do not understand it.. i shall re-explain

OpenStudy (goformit100):

sure please

OpenStudy (perl):

so the answer is 10

OpenStudy (hexagon001):

the first example is 1+2=5.. the second term is 2.. square this number..this gives 4 now add the first term, which is 1, to 4. 1+4 =5 the second example follows the same principle

OpenStudy (hexagon001):

3+4 =19 take the secon term, which is 4 now square this term. 4^2=16 now add 16 to the first term in the expression(3) so 3+16=19